OpenStudy (anonymous):

∫_2^(-2) {(3dt/(4+3t^2))}

OpenStudy (anonymous):

http://integrals.wolfram.com/index.jsp

OpenStudy (anonymous):

\[\int\limits_{-2}^{2}{\frac{3dt}{4+3t^2}}\]

OpenStudy (anonymous):

i dont want to know the answer i need to know how to do it

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

here we go: \[1/4 * 3/ (4/4 +(\sqrt{3}x/2)^{2}=3/4 * 1/(1+(\sqrt{3}x/2)^{2}\]

OpenStudy (anonymous):

this is under integral

OpenStudy (anonymous):

I'm feeling an arctan flavor here:\[\frac{3dt}{4+3t^2}=\frac{3dt}{(4)(1+\frac{3t^2}{4})}=\frac{3}{4}\cdot\frac{dt}{1+(\frac{\sqrt{3}t}{2})^2}\]Move the 3/4 to the outside of the integral, "balance" the squared portion, use arctan...

OpenStudy (anonymous):

now you have: \[3/2\int\limits_{ }^{}d(\sqrt{3}x/2)/(1+(\sqrt{3}x/2)^{2} *2/\sqrt{3}=\sqrt{3}/2 \tan ^{-1}(\sqrt{3}x/2)\]

OpenStudy (anonymous):

put your (-2,2) & calculate

OpenStudy (anonymous):

are you OK with this? generally it's substitution u=sq.rt(3) x/2

OpenStudy (anonymous):

its not really clear what u did but i'll write it out thank you inik and londovir

OpenStudy (anonymous):

The main thing is if you ever see an integral with no variable in the numerator (perhaps just a constant number), but with a squared variable in the denominator, it's often inverse trig...when it's a + in the denominator, think inverse tangent (or arctan)...

OpenStudy (anonymous):

or we didnt learn this yet...okay thank you